OpenStudy (anonymous):

The expression cos 4x cos 3x + sin 4x sin 3x is equivalent to

OpenStudy (zzr0ck3r):

cos(a-b)=cos(a)cos(b)+sin(a)sin(b)

OpenStudy (zzr0ck3r):

you have a = 4x and b = 3x

OpenStudy (zzr0ck3r):

note the order does not matter because cos(-x) = cos(x)